OpenStudy (anonymous):

d/dx (xe^(lnx^2)) =?

OpenStudy (anonymous):

since e^x and lnZ(x) are inverse operations, it is true that:\[\ln(e^x) = x\] and\[e^{\ln(x)} = x\] using same logic:\[e^{\ln(x^2)} = x^2\] we aer now looking for d/dx (x*x^2) = d/dx x^3 i think you can handle it now ^_^

OpenStudy (anonymous):

x^3?

OpenStudy (anonymous):

x*e^(lnx^2) = x^3 so you have to derive x^3

OpenStudy (anonymous):

i understand now :)

OpenStudy (anonymous):

3x^2 thanks :)